OpenStudy (anonymous):

Find the difference quotient of f, that is, find f(x+h) - f(x) --------- h for these functions: f(x) = -4x+1 f(x) = x^2 -x +4

OpenStudy (bahrom7893):

1) -4 2) 2x-1

OpenStudy (anonymous):

I originally put 2x-1 & it was not the right answer

OpenStudy (bahrom7893):

is it the limit as h->0?

OpenStudy (anonymous):

h does not equal 0

OpenStudy (anonymous):

yeah, 2x-1 is the derivative, but its not asking for the lim as h-->0

OpenStudy (bahrom7893):

ohh I see wait up..for the 2nd one Here: (x+h)^2-(x+h)+4 - (x^2 -x +4) --------------------------- h

OpenStudy (anonymous):

\[f(x+h) = (x+h)^{2}-(x+h)+4\] \[= x^{2}+2xh+h^{2}-x-h+4\] \[\frac{f(x+h)-f(x)}{h} = \frac{x^{2}+2xh+h^{2}-x-h+4-x^{2}+x-4}{h} = \frac{2xh+h^{2}-h}{h}\]

OpenStudy (bahrom7893):

Simplify: [x^2+2xh+h^2-x-h+4-x^2-x+4]/h

OpenStudy (anonymous):

Dividing out the h we get: \[2x+h-1\] which would be you answer

OpenStudy (anonymous):

notice if you took the lim as h went to 0 you would get the derivative, which is 2x-1

OpenStudy (anonymous):

thanks!